OpenStudy (anonymous):

find the LMC of 18 and 27

OpenStudy (anonymous):

which is the least common multiple

OpenStudy (igreen):

Least Common Multiple here is the smallest number that can get into both 18 and 27 which doesn't include the number 1.

OpenStudy (igreen):

Let's try 2. 18 / 2 = 9. So 18 is a multiple of 2. 27 / 2 = 13.5 So 27 is not a multiple of 2. Since 2 cannot get into 27, it is not the LCM of 18 and 27.

OpenStudy (igreen):

Let's try 3. 18 / 3 = 6 So 18 is a multiple of 3 27 / 3 = 9 So 27 is a multiple of 3 Since 3 can get into both 18 and 27, 3 is the LCM of 18 and 27.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @Tttttttcoockiemonster

OpenStudy (anonymous):

Thanx so much!!!